# One Wave to Explain Them All: A Unifying Perspective on Post-hoc Explainability

### v1

@@ Abstract
Despite the growing use of deep neural networks in safety-critical decision-making, their inherent black-box nature hinders transparency and interpretability. Explainable AI (XAI) methods have thus emerged to understand a model’s internal workings, and notably attribution methods also called Saliency maps. Conventional attribution methods typically identify the locations – the where – of significant regions within an input. However, because they overlook the inherent structure of the input data, these methods often fail to interpret what these regions represent in terms of structural components (e.g., textures in images or transients in sounds). Furthermore, existing methods are usually tailored to a single data modality, limiting their generalizability. In this paper, we propose leveraging the wavelet domain as a robust mathematical foundation for attribution. Our approach, the Wavelet Attribution Method (WAM), extends the existing gradient-based feature attributions into the wavelet domain, providing a unified framework for explaining classifiers across images, audio, and 3D shapes. Empirical evaluations demonstrate that WAM matches or surpasses state-of-the-art methods across faithfulness metrics and models in image, audio, and 3D explainability. Finally, we show how our method explains not only the where – the important parts of the input – but also the what – the relevant patterns in terms of structural components.
11footnotetext: *Alphabetical ordering.
@@ 1 Introduction
Deep neural networks are increasingly being deployed in various applications, such as medicine, transportation, robotics, or finance (Pooch et al., 2020; Sun et al., 2022; Redmon et al., 2016; Thimonier et al., 2024). These networks often make critical decisions, such as detecting tumors in medical images or identifying obstacles in autonomous driving, yet the underlying decision-making process is difficult to interpret due to the black-box nature of the models.
This opacity has motivated the rise of explainable AI (XAI) techniques to provide human-understandable explanations for model decisions. While XAI has been predominantly applied in image classification, it is also extending into other fields, such as audio and 3D shape classification (Parekh, 2023; Paissan et al., 2024; Chen et al., 2021; Zheng et al., 2019).
Among these techniques, feature attribution methods – specifically gradient-based methods for generating saliency maps (heatmaps that highlight important input features, Zeiler & Fergus, 2014) are prevalent. These gradient-based methods (Shrikumar et al., 2017; Sundararajan et al., 2017; Smilkov et al., 2017) considered as efficient and reliable for interpreting model behavior (Crabbé & van der Schaar, 2023; Wang & Wang, 2021; Xue et al., 2023).
Feature attribution involves decomposing a model’s decision within a specific “explanation” domain. Traditionally, saliency mapping relied on the pixel domain as this domain. However, pixel-based explanations flatten the hierarchical and spatial relationships inherent in images, effectively collapsing their structural properties. In addition, the pixel domain is only relevant when the input modality is an image. Instead, decomposing the model’s decision in the wavelet domain, which preserves the inter-scale dependencies of an input modality, could enable saliency-based methods to account for the image structure in the explanation. Besides, the wavelet domain is defined for any input dimension (images being an input of dimension 2), thus enabling a natural generalization of saliency mapping to modalities such as audio or 3D volumes.
This work introduces the Wavelet Attribution Method (WAM), a universal feature attribution method. By unifying and extending existing methods, notably SmoothGrad (Smilkov et al., 2017) and Integrated Gradients (Sundararajan et al., 2017) within the wavelet domain, we enable their application to any modality defined over a continuous space, moving beyond the limitations of the pixel domain. As illustrated in Figure 1, our approach involves computing the gradient of a classification model’s prediction with respect to the wavelet decomposition of the input signal. We then produce smooth explanations by either averaging over noisy inputs or integrating along the prediction path.
Operating in the wavelet domain, WAM isolates the contribution of the different scales within the input signals to the model’s prediction, providing deeper insights into a model’s decision-making process. We illustrate these insights by revisiting the meaningful perturbation framework withWAM, or by carrying out noise and overlapping experiments on audio samples. Quantitative evaluations demonstrate that WAM outperforms existing attribution methods across a range of topologies, modalities, and metrics, underscoring its utility in addressing critical challenges in image, audio, and 3D shape classification.
Figure 1: Explaining any modality by decomposing the model’s decision in the wavelet domain. By computing the gradient of the model’s prediction in the wavelet domain, we provide a unified approach to explain the decision of audio, image, and shape classifiers that preserves the structural properties of the input modalities.
@@ 2 Related works
Images explainability.
Computer vision has supported the development of numerous post-hoc explainability methods (Baehrens et al., 2010), with attribution methods being the most popular. Post-hoc methods are applied on trained model and aim to rank, i.e., estimate an importance for each pixel or region of an image, reflecting its contribution to the score predicted by the model. Many methods have been proposed, which can be classified into two categories: White-box methods, which require access to gradients, and Black-box methods, which use perturbations on the input space.
White-box approaches leverage internal model information, such as gradients, to explain predictions. saliency maps were first introduced by Baehrens et al. (2010) and later refined in Simonyan et al. (2014); Zeiler & Fergus (2014); Springenberg et al. (2014); Sundararajan et al. (2017); Smilkov et al. (2017); or Muzellec et al. (2023). These methods calculate the gradient of the classification score with respect to the input pixels using backpropagation. However, gradients can be noisy in large vision models (Smilkov et al., 2017) and may yield misleading importance estimates due to their focus on infinitesimal input variations (Ghalebikesabi et al., 2021).On the other hand, Black-box methods rely on input perturbations without requiring access to internal model parameters. Techniques like Occlusion (Zeiler & Fergus, 2014), LIME (Ribeiro et al., 2016), RISE (Petsiuk et al., 2018), Sobol (Fel et al., 2021), and HSIC (Novello et al., 2022) generate importance maps by observing changes in the classification score when the input image is altered. For example, Occlusion uses binary masks to systematically occlude regions of the image, while RISE and HSIC apply random masks to perturb multiple regions simultaneously.
However, traditional attribution methods focus on where important regions in the image are. However, they fail to address what these regions represent in terms of meaningful, higher-level concepts – a gap that more recent research is attempting to fill (Kim et al., 2018; Ghorbani et al., 2019; Fel et al., 2023b; Zhang et al., 2021; Graziani et al., 2023; Fel et al., 2023b; a). Moreover, these methods often under-utilize the inherent structure of images, (recent work are beginning to explore attribution in the frequency domain, e.g. Muzellec et al., 2023).
Audio explainability.
@@
Previous work on post-hoc audio explainability methods has mainly expanded in three directions. The first explored the use of saliency methods to highlight key features for audio classifiers processing spectrograms (Becker et al., 2024; Won et al., 2019) or 1D waveforms (Muckenhirn et al., 2019). Moreover, while the use of time-frequency representations for classification and explanations is frequent in this regard, wavelet representations have not been explored previously for explanations. The second direction involves variants of LIME (Ribeiro et al., 2016) algorithm, proposed for different types of audio classification tasks (Mishra et al., 2017; 2020; Haunschmid et al., 2020; Chowdhury et al., 2021; Wullenweber et al., 2022). The third has pursued the development of methods to generate listenable interpretations for audio classifiers by leveraging the hidden representations (Parekh et al., 2022; Paissan et al., 2024). LIME-based methods suffer from the issue of high computational costs of explanations due to a large number of forward passes per sample. Recent methods for listenable interpretations require access to hidden layers and train separate modules and are thus unsuitable as post-hoc explainers.
3D explainability.
@@
3D data generally comes in two main formats: point clouds and voxels. Point clouds offer an exact representation of the data but are unstructured. Voxels, conversely, are a discretized but structured representation of the data, making them suitable for processing with techniques such as 3D convolutions. Most explainability techniques for 3D data focused on explaining point clouds. Chen et al. (2021); Schinagl et al. (2022); Gupta et al. (2020), and Zheng et al. (2019) introduced techniques to generate visual explanations for interpretability of 3D object detection and classification networks. They highlight critical features in point cloud data by adapting 2D image-based saliency techniques (Gupta et al., 2020; Zheng et al., 2019), by using a perturbation-based approach (Schinagl et al., 2022) or by proposing a 3D variant of LIME (Tan & Kotthaus, 2022). Explainability on 3D volumes remains limited. A few works (Yang et al., 2018; Mamalakis et al., 2023; Gotkowski et al., 2021) have proposed attention maps on 2D slices of 3D medical scans using 3D-GradCAM.
3D and 1D explainability techniques often reproject the model’s decision onto a 2D-pixel domain. However, this projection filters out the intrinsic properties of the original signal, such as its temporal or spatial depth, resulting in an incomplete representation. This process, therefore, constitutes an improper way of generalizing attribution methods, as it disregards essential features of the original signal’s structure. In addition, the pixel domain itself is limited for explainability. More broadly, we note that the literature has only recently started discussing the expressiveness of the “explanation” domain, e.g., through the lenses of concepts, and still overlooks the broader applicability across modalities. This work contributes to the ongoing discussion by evaluating how the wavelet domain can simultaneously address these concerns.
3 Methods
Notations & Background.
@@ 3 Methods
Throughout, we let $\mathcal{X}=(\Omega,\mathcal{F},\mu)$ be a measure space with set $\Omega$, $\sigma$-algebra $\mathcal{F}$, and measure $\mu$. We denote by $\mathcal{H}=\mathbb{L}^{2}(\mathcal{X},\mu)$ the Hilbert space of square-integrable functions on $\mathcal{X}$. Let $\bm{f}\in\mathcal{H}$ represent a predictor function (e.g., a classifier), which maps an input $\bm{x}\in\mathcal{X}$ to an output $\bm{f}(\bm{x})\in\mathcal{Y}$. We denote $\bm{g}\in\mathcal{H}$ a generic, square-integrable function.
A wavelet is an integrable function $\bm{\psi}\in\mathcal{H}$ that is normalized, centered at 0, and has zero average (i.e., $\int\bm{\psi}(\bm{x})\,\mathrm{d}\bm{x}=0$). Unlike a sine wave, a wavelet is localized in both space and frequency domains. This localization allows dilations of the wavelet to analyze different frequency intervals (scales) while translations enable analysis at different spatial locations. To compute an image’s continuous wavelet transform (CWT), we first define a filter bank $\mathcal{D}$ derived from the original wavelet $\bm{\psi}$, using a scale factor $\bm{\lambda}>0$ and 2D translation $\bm{b}$. The filter bank $\mathcal{D}$ is given by
$\mathcal{D}=\left\{\bm{\psi}_{\bm{\lambda},\bm{b}}(\bm{x})=\frac{1}{\sqrt{\bm{\lambda}}}\bm{\psi}\left(\frac{\bm{x}-\bm{b}}{\bm{\lambda}}\right)\right\}_{\bm{b}\in\mathbb{R}^{2},\,\bm{\lambda}>0}.$
The continuous wavelet transform of a function $\bm{g}\in\mathcal{H}$ at scale $\bm{\lambda}$ and location $\bm{x}$ is given by
$\mathcal{W}(\bm{g})(\bm{\lambda},\bm{x})=\int_{-\infty}^{+\infty}\bm{g}(\bm{b})\frac{1}{\sqrt{\bm{\lambda}}}\bm{\psi}^{*}\left(\frac{\bm{b}-\bm{x}}{\bm{\lambda}}\right)\mathrm{d}\bm{b},$

### v2

# One Wave To Explain Them All: A Unifying Perspective On Feature Attribution

@@ Abstract
Feature attribution methods aim to improve the transparency of deep neural networks by identifying the input features that influence a model’s decision. Pixel-based heatmaps have become the standard for attributing features to high-dimensional inputs, such as images, audio representations, and volumes. While intuitive and convenient, these pixel-based attributions fail to capture the underlying structure of the data. Moreover, the choice of domain for computing attributions has often been overlooked. This work demonstrates that the wavelet domain allows for informative and meaningful attributions. It handles any input dimension and offers a unified approach to feature attribution. Our method, the Wavelet Attribution Method (WAM), leverages the spatial and scale-localized properties of wavelet coefficients to provide explanations that capture both the where and what of a model’s decision-making process. We show that WAM quantitatively matches or outperforms existing gradient-based methods across multiple modalities, including audio, images, and volumes. Additionally, we discuss how WAM bridges attribution with broader aspects of model robustness and transparency. Project page: https://gabrielkasmi.github.io/wam/.
Interpretability, Feature attribution, Wavelets, Images, Audio, 3D, Volumes
@@ 1 Introduction
Figure 1: Explaining any input modality by decomposing the model’s decision in the wavelet domain. WAM computes the gradient of the model’s prediction with respect to the wavelet coefficients of the input modality (audio, image, volume). Unlike pixels, wavelet coefficients preserve structural information about the input signal, offering deeper insights into the model’s behavior and going beyond where it focuses.
Deep neural networks are increasingly being deployed in various applications. However, their opacity poses a significant challenge, especially in safety-critical domains, where understanding the decision-making process is crucial, for instance, in tumor detection (Pooch et al., 2020) or obstacle identification (Sun et al., 2022).
This opacity motivated the rise of Explainable Artificial Intelligence (XAI) techniques to provide human-understandable explanations of model decisions. While XAI has been predominantly applied in image classification, it is also extending into other fields, such as audio and volumes classification (Parekh, 2023; Paissan et al., 2024; Chen et al., 2021; Zheng et al., 2019). Currently, the most popular XAI methods are feature attribution methods (Shrikumar et al., 2017; Sundararajan et al., 2017; Smilkov et al., 2017; Fel et al., 2021; Novello et al., 2022; Muzellec et al., 2024) which consist in generating saliency maps, i.e., heatmaps highlighting the important pixels on the input image (Zeiler & Fergus, 2014b). The same principle has been expanded for audio and volumes, where most feature attributions are computed on 2D projections of the input modality. However, the pixel domain does not inherently capture the structural properties of signals. So far, existing works have overlooked the question of the domain in which feature attribution is computed. However, signal processing theory states that meaningful features often exist at multiple scales and are better represented in transformed domains that account for spatial and frequency information. This underlines the need for an alternative representations that preserves these properties to provide more faithful explanations.
Wavelets offer a hierarchical decomposition that retains both spatial and frequency information, unlike pixel-based methods that lose structural context. This makes wavelets a stronger foundation for interpreting model decisions across diverse modalities (or signals), as wavelets are inherently low-level features defined across various signal dimensions. This work introduces the Wavelet Attribution Method (WAM), which leverages wavelet coefficients as the basis features for computing attributions. Therefore, WAM provides a unification of the feature attribution domain on components (wavelet coefficients) that are interpretable as low-level features of the input signal. We expand popular feature attribution methods, namely SmoothGrad (Smilkov et al., 2017) and Integrated Gradients (Sundararajan et al., 2017), within the wavelet domain, thus providing a generalization of feature attribution to any square-integrable modality. WAM preserves the original structure of the data without converting it into a 2D representation.
Figure 1 illustrates our approach, which computes the gradient of a classification model with respect to the wavelet coefficients of the input signal. These coefficients are defined for any square-integrable function, regardless of dimension, and are localized in both space and scale, allowing them to capture information across different scales. For images and volumes, wavelet coefficients capture shapes, edges, and textures, while for audio signals, they represent transient patterns. Consequently, feature attribution in the wavelet domain enables us to interpret audio, highlight important areas and patterns in images, or identify key regions in volumes, all in a post-hoc manner, i.e., using only a trained model and a single backward pass.
We demonstrate the superiority of WAM through extensive empirical evaluations across a diverse set of metrics, model architectures, and datasets, showing its robustness and versatility. We also discuss the novel insights and connections that our method permits, such as connecting feature attribution and the characterization of a model’s robustness or filtering the important part of an audio signal without requiring a to train an explanation method.
In summary, the following are the key contributions of this work:
  • We introduce WAM, a novel feature attribution method that operates in the wavelet domain, ensuring that the attributions align with the intrinsic properties of the data, whether it be audio, images, or volumes.
  • We demonstrate the effectiveness of WAM through extensive empirical evaluations across diverse models, datasets, and performance metrics.
  • We illustrate the relevance of transitioning to the wavelet domain for feature attribution by showcasing how WAM offers novel insights into model transparency.
@@ 2 Related works
2.1 Feature attribution methods
Images.
Computer vision has supported the development of numerous post-hoc feature attribution methods (Baehrens et al., 2010). These techniques are applied to a trained model and estimate the importance of each pixel or region of an image based on its contribution to the model’s prediction, i.e., estimate a saliency map. Post-hoc methods either leverage internal model information, such as gradients (Baehrens et al., 2010; Simonyan et al., 2014; Zeiler & Fergus, 2014b; Springenberg et al., 2014; Sundararajan et al., 2017; Smilkov et al., 2017; Muzellec et al., 2024; Han et al., 2024; Binder et al., 2016) or apply perturbations to the input space (Lundberg & Lee, 2017; Ribeiro et al., 2016; Petsiuk et al., 2018; Fel et al., 2021; Novello et al., 2022; Kasmi et al., 2023a).
Audio.
@@ 2.1 Feature attribution methods
The methods introduced in computer vision have been extended to audio classification in three directions. The first one explores the use of saliency methods to highlight key features for audio classifiers processing spectrograms (Becker et al., 2024; Won et al., 2019) or waveforms (Muckenhirn et al., 2019). The second direction involves variants of LIME (Ribeiro et al., 2016) algorithm, proposed for different types of audio classification tasks (Mishra et al., 2017, 2020; Haunschmid et al., 2020; Chowdhury et al., 2021; Wullenweber et al., 2022). The third pursues the development of methods to generate listenable interpretations for audio classifiers by leveraging the hidden representations (Parekh et al., 2022; Paissan et al., 2024).
Volumes.
@@
3D data can be represented as point clouds or voxels. Point clouds offer an exact representation of the data but are unstructured. Voxels, conversely, are a discretized but structured representation of the data, making them suitable for processing with techniques such as 3D convolutions. Most explainability techniques for 3D data focused on explaining point clouds. Chen et al. (2021); Schinagl et al. (2022); Gupta et al. (2020); Zheng et al. (2019) introduced techniques to generate visual explanations for interpretability of 3D object detection and classification networks. They highlight critical features in point cloud by adapting 2D image-based saliency techniques (Gupta et al., 2020; Zheng et al., 2019), by using a perturbation-based approach (Schinagl et al., 2022) or by proposing a 3D variant of LIME (Tan & Kotthaus, 2022). Explainability on volumes remains limited. A few works (Yang et al., 2018; Mamalakis et al., 2023; Gotkowski et al., 2021) have proposed attention maps on 2D slices of 3D medical scans using 3D-GradCAM.
Current feature attribution methods rely on a feature space that overlooks the inherent temporal, spatial, or geometric relationships within the data. Even worse, projecting explanations into the two-dimensional pixel domain for 1D audio or 3D volumes further distorts these relationships, resulting in explanations that fail to capture the full complexity of the model’s decision-making process, ultimately reducing the relevance of the attributions.
2.2 Wavelet-based explainability
From an explainability perspective, we argue that the wavelet domain is more informative for feature attribution than the pixel domain, as it enables the extraction of interpretable low-level features from input signals. In particular, the dyadic wavelet decomposition of a signal isolates distinct patterns within the input signal. In images, wavelet coefficients capture edges, textures, and patterns at different scales and orientations, aligning with human perception. For instance, wavelets coincide in some cases with Gabor filters (Gabor, 1946), which have been widely used for pattern analysis. In audio, they separate transient components from slowly varying structures, facilitating the identification of speech formats or sudden bursts of sound.
Surprisingly, hardly a handful of works explored the use of wavelets for post-hoc explainability. CartoonX (Kolek et al., 2022) and ShearletX (Kolek et al., 2023) extended the Meaningful Perturbation framework (Fong & Vedaldi, 2017) into the wavelet and shearlet domains, respectively, allowing to highlight not just salient pixel regions but also relevant textures and edge structures. Recently, the Wavelet Scale Attribution Method (WCAM, Kasmi et al., 2023a), transposed Sobol-based feature attribution (Fel et al., 2021) to the wavelet domain, demonstrating the feasibility of multiscale attributions for images. Wavelet coefficients have been used to construct scattering transforms (Bruna & Mallat, 2013; Andén & Mallat, 2014), i.e., fixed feature extractors that create a translation-invariant representation of an input, stable to deformations. This aligns with the properties of trained feature extractors like convolutional neural networks (CNNs) without any training. This model and its descendants, such as the scattering spectra (Cheng et al., 2024), can be considered intrinsically interpretable models (Flora et al., 2022) and have been applied in fields where model understanding is crucial, such as finance (Leonarduzzi et al., 2019) and astrophysics (Cheng & Ménard, 2021).
3 Feature attribution in the wavelet domain
Figure 2: WAM for images. Our method decomposes important components at multiple scales, revealing what the model focuses on. The same spatial location encodes different structural elements at different scales: the model predicts the elephant because of its trunk and ears but needs both the fine texture details (blue), the intermediate-scale contours (red), and the coarse edges (yellow). Such a decomposition is not possible with traditional feature attribution in the pixel domain, and this hierarchical decomposition provides an interpretable view of model attributions.
Notations.
@@ 3 Feature attribution in the wavelet domain
Throughout, we let $\mathcal{X}=(\Omega,\mathcal{F},\mu)$ be a measure space with set $\Omega$, $\sigma$-algebra $\mathcal{F}$, and measure $\mu$. $\mathcal{H}=L^{2}(\mathcal{X},\mu)$ denotes the Hilbert space of square-integrable functions on $\mathcal{X}$. Let $\bm{f}\in\mathcal{H}$ represent a predictor function (e.g., a classifier), which maps an input $\bm{x}\in\mathcal{X}$ to an output $\bm{f}(\bm{x})\in\mathcal{Y}$. We denote $\bm{g}\in\mathcal{H}$ a square-integrable function.
3.1 Multiscale decompositions
Overview.
Wavelets decompose signals into components that retain both spatial and frequency information, unlike Fourier transforms, which capture only frequency and discard spatial or temporal context. The choice of the wavelet determines what kind of structural components will be highlighted on the image. Its choice is therefore grounded in the application at hand. We refer the reader to Section A.2 for a discussion of the choice of the mother wavelet and its impact on the interpretation of the WAM.
Definition.
A wavelet is an integrable function $\bm{\psi}\in\mathcal{H}=L^{2}(\mathbb{R}^{n})$ that is normalized, centered at the origin, and has zero mean (i.e., $\int_{\mathbb{R}^{n}}\bm{\psi}(\bm{x})\,\mathrm{d}\bm{x}=0$). In contrast to sine waves, which are localized in frequency but not in space, wavelets are localized in both space and frequency domains. To analyze a function or a signal $\bm{g}\in\mathcal{H}$, we define a family of functions $\mathcal{D}$ obtained by dilating and translating a mother wavelet $\bm{\psi}$:
$\mathcal{D}=\left\{\bm{\psi}_{\bm{\lambda},\bm{b}}(\bm{x})=\frac{1}{\bm{\lambda}^{n/2}}\bm{\psi}\left(\frac{\bm{x}-\bm{b}}{\bm{\lambda}}\right)\right\}_{\bm{b}\in\mathbb{R}^{n},\,\bm{\lambda}>0},$ (1)
where $\bm{b}$ is a translation and $\bm{\lambda}$ is a scale factor. This family gives rise to the continuous wavelet transform (CWT) of $\bm{g}$, defined as
$\mathcal{W}(\bm{g})(\bm{\lambda},\bm{b})$ $=\langle\bm{g},\bm{\psi}_{\bm{\lambda},\bm{b}}\rangle$
$=\int_{\mathbb{R}^{n}}\bm{g}(\bm{x})\frac{1}{\bm{\lambda}^{n/2}}\overline{\bm{\psi}}\left(\frac{\bm{x}-\bm{b}}{\bm{\lambda}}\right)\mathrm{d}\bm{x}.$ (2)
This operation can be interpreted as a convolution with a time-reversed and dilated version of the complex-conjugate $\overline{\bm{\psi}}$ of $\bm{\psi}$ (i.e., a linear filtering operation), and thus defines a linear operator on $\mathcal{H}$ (Mallat, 2008).
Discrete Wavelet Transform.
When the scale parameter is restricted to dyadic values $\bm{\Lambda}=\{2^{j}:j\in\mathbb{Z}\}$ and translations lie on a discrete grid, the continuous family becomes a countable set. Under appropriate admissibility and regularity conditions on $\bm{\psi}$, this discrete family forms an orthonormal basis or a tight frame for $\mathcal{H}$, giving rise to the Discrete Wavelet Transform (DWT).
The DWT is defined by projecting a signal $\bm{g}\in L^{2}(\mathbb{R}^{n})$ onto a dyadic family of wavelets obtained by discretizing the scale and translation parameters. For dyadic scales $\bm{\lambda}=2^{j}$ and integer translations $\bm{b}=2^{j}k$, we define the family
$\left\{\bm{\psi}_{j,k}(\bm{x})=2^{jn/2}\bm{\psi}(2^{j}\bm{x}-k)\right\}_{j,k\in\mathbb{Z}}.$ (3)
The DWT coefficients are given by the inner products
$\mathcal{W}_{\text{DWT}}(\bm{g})(j,k)=\langle\bm{g},\bm{\psi}_{j,k}\rangle=\int_{\mathbb{R}^{n}}\bm{g}(\bm{x})\overline{\bm{\psi}_{j,k}(\bm{x})}\,\mathrm{d}\bm{x}.$ (4)
These coefficients represent the content of $\bm{g}$ at scale $2^{-j}$ and position $2^{-j}k$. Under the appropriate admissibility and regularity conditions on $\bm{\psi}$, the DWT is an invertible linear transform: the signal $\bm{g}$ can be exactly reconstructed from its wavelet coefficients $\mathcal{W}_{\text{DWT}}(\bm{g})(j,k)$ via the inverse transform
$\bm{g}(\bm{x})=\sum_{j\in\mathbb{Z}}\sum_{k\in\mathbb{Z}}\mathcal{W}_{\text{DWT}}(\bm{g})(j,k)\,\bm{\psi}_{j,k}(\bm{x}).$ (5)